Question: Distance between C and A 353 A BC 49 АВС a bc a BC A bc 352 a bC 46 a B c 98 Please answer the following question(s): 1. Based on the data provided, what is the distance between genes C and A? b. 0.1 cM C.1 cM d. 10 cM e. 100 cM Safety inthe Lab doc G) how all aety,in the.labdoc
Answer #1

To find out the distance between genes C and A, we have to calculate the recombination frequency between them. For this we have to take the double cross overs ( having lowest number of indivduals) along with the single cross overs between C and A.

The SCO individuals are Cab - 46 and cAB - 49 and the DCO individuals are CaB - 2 and cAb - 3

Total number of individuals are 353 + 352 + 49 + 46 + 92 + 98 + 2 + 3 = 1000

So, cross over frequency between A and C is (46+49 + 2 + 3 / 1000) x 100 = (100/ 1000) x 100

= 10 cM